OpenStudy (anonymous):

a baseball player scored 14 runs in 18 games. this season the baseball player scored 12 runs in 15 games. find the number of runs scored per game in each season. round to the nearest hundreth. then identify the season in which the player scored more runs per game.

OpenStudy (ghazi):

in 18 games 14 runs were made therefore in 1 game number of runs = 14/18=7/9=0.78 now, in 15 games 12 runs were made therefore in one game runs made= 12/15= 4/5=0.8 in second session he scored more runs

OpenStudy (anonymous):

thank u xD the second session is 0.78?

OpenStudy (ghazi):

:) well the second session is 12/15= 0.8

OpenStudy (ghazi):

@pudding is it clear?

OpenStudy (anonymous):

why is 0.8 higher than 0.78??

OpenStudy (ghazi):

well 0.8 means 0.80 and it is greater than 0.78...you've to round to the nearest hundred too

OpenStudy (anonymous):

ohh i see (: thank u so much

OpenStudy (ghazi):

0.80 is 0.02 more than 0.78.... 0.78+0.02= 0.80

OpenStudy (ghazi):

:) YW

OpenStudy (anonymous):

could u help me with one more pls?

OpenStudy (ghazi):

sure

OpenStudy (anonymous):

you are getting ready to make a 640 mile car trip. in general, your car can drive 160 miles on 5 gallons of gasoline. how many gallons of gasoline will you use for your trip? you started out with 4 gallons of gasoline in your car and gasoline is $2.29 per gallon.how much will you spend on gasoline on the trip?

OpenStudy (ghazi):

160 miles in 5 gallons 1 mile in 5/160 gallons = 0.03125 gallons therefore for 640 mile ,number of gallon = 640*0.03125=20 gallons i already had 4 gallons so i need 16 more 1 gallon cost 2.29 therefore 16 = 16*2.29= $ 36.64 :)))

OpenStudy (ghazi):

@pudding is this cool ?

OpenStudy (anonymous):

@ghazi wait im tryin 2 read it and then makin it on my own :p

OpenStudy (anonymous):

if i have 4 gallons why do i need 16 more? where does it say 16?

OpenStudy (anonymous):

i dont get that part

OpenStudy (ghazi):

well!! we need 20 gallons to travel 640 miles and we started journey with 4 gallons so we need 16 more to go 640 miles

OpenStudy (anonymous):

ohh so 20-4

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

thank you ((; that's all for today hugs

OpenStudy (ghazi):

hugs to you too ...you're cool !!

OpenStudy (anonymous):

awhh thts so sweet of u!! ur like so awesome xD

OpenStudy (ghazi):

xD

OpenStudy (anonymous):

you just answered my math homework , thank you soooo much